# Towards Self‑Evolving Benchmarks: Synthesizing Agent Trajectories via Test‑Time Exploration under Validate‑by‑Reproduce Paradigm

### v1

@@ Abstract
Recent advances in large language models (LLMs) and agent system designs have empowered agents with unprecedented levels of capability. However, existing agent benchmarks are showing a trend of rapid ceiling-hitting by newly developed agents, making it difficult to meet the demands for evaluating agent abilities. To address this problem, we propose the Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution (TRACE) framework. This framework takes an original task from an existing benchmark and encourages agents to freely explore and evolve it into a new task with higher difficulty while recording validatable agent trajectories. The framework proceeds in three stages: (1) evolutionary proposal mining, which provides task evolution proposals through preliminary exploration and divergent thinking; (2) problem formation and free exploration, where proposals are conceptualized into feasible problem candidates and the agents then explore them freely while recording their execution trajectories; and (3) multi-level validation, which ensures that the evolved tasks are accompanied by validatable and reproducible trajectories. Experiments on the GAIA benchmark demonstrate that the TRACE framework consistently enhances task complexity while improving the reliability of correctness through validatable execution trajectories. This work marks a paradigm shift from static, manually curated benchmarks to dynamic, self-evolving evaluation systems, providing a sustainable and challenging runway for agent development.
@@ 5.2 Experimental Results and Analysis
Decreased Performance on New Tasks.
@@
From Seed to Spark.
@@
Table 2: Model Evaluation Results Pass@1

### v2

@@ Abstract
Recent advances in large language models (LLMs) and agent system designs have empowered agents with unprecedented levels of capability. However, existing agent benchmarks are showing a trend of rapid ceiling-hitting by newly developed agents, making it difficult to meet the demands for evaluating agent abilities. To address this problem, we propose the Trajectory-based Validated-by-Reproducing Agent-benchmark Complexity Evolution (TRACE) framework. This framework takes an original task from an existing benchmark and encourages agents to freely explore and evolve it into a new task with higher difficulty while recording validatable agent trajectories. The framework proceeds in three stages: (1) evolutionary proposal mining, which provides task evolution proposals through preliminary exploration and divergent thinking; (2) problem formation and free exploration, where proposals are conceptualized into feasible problem candidates and the agents then explore them freely while recording their execution trajectories; and (3) multi-level validation, which ensures that the evolved tasks are accompanied by validatable and reproducible trajectories. Experiments on the GAIA benchmark demonstrate that the TRACE framework consistently enhances task complexity while improving the reliability of correctness through validatable execution trajectories. In addition, our framework can successfully adapt to and improve reasoning datasets represented by AIME-2024. This work marks a paradigm shift from static, manually curated benchmarks to dynamic, self-evolving evaluation systems, providing a sustainable and challenging runway for agent development.
@@ 5.2 Experimental Results and Analysis
5.2.1 Decreased Performance on New Tasks.
To assess the generality of our evolutionary framework, we compare model performance before and after evolution on two benchmarks: GAIA and AIME-2024. In both cases, models exhibit clear performance degradation on the evolved tasks.
GAIA
@@
AIME-2024
Table 3 presents the performance of three reasoning models(DeepSeek-R1-Distill-Qwen-7B, DeepSeek-R1-Distill-Qwen-32B, Qwen3-235B-A22B) of different scales on the AIME-2024 benchmark, as well as its second and fourth evolutionary benchmarks. Average Acc. refers to the average accuracy over 10 test runs on the benchmark, while Average Length denotes the average number of tokens in the chain-of-thought generated by the model. We can observe that after four rounds of evolution, the model’s performance on the benchmark has significantly declined, and the token count has noticeably increased. For instance, the Average Acc. of the Qwen3-235B-A22B model decreased by 22.33%, and the average token count increased by over 8000. These experimental data demonstrate the significant change in the difficulty of the benchmark questions.
5.2.2 From Seed to Spark.
@@
Table 2: Model Evaluation Results Pass@1 on GAIA benchmark
@@
[…]
Table 3: Model Evaluation Results on AIME-2024 benchmark
Models DeepSeek-R1-Distill-Qwen-7B DeepSeek-R1-Distill-Qwen-32B Qwen3-235B-A22B
[…]
ROUND 2
[…]
ROUND 4
@@ B.2 Task Evolution Demonstrations
Case 9 Original Task: Jen enters a lottery by picking 44 distinct numbers from $S=\{1,2,3,\cdots,9,10\}.$ 44 numbers are randomly chosen from $S.$ She wins a prize if at least two of her numbers were 22 of the randomly chosen numbers, and wins the grand prize if all four of her numbers were the randomly chosen numbers. The probability of her winning the grand prize given that she won a prize is $\tfrac{m}{n}$ where $m$ and $n$ are relatively prime positive integers. Find $m+n$. Evolved Task: Jen picks a set $J$ of 4 distinct numbers from $S=\{1,2,\dots,10\}$. Tom then picks a set $T$ of 4 distinct numbers from $S$, chosen uniformly at random from all such sets that have exactly one number in common with $J$ (i.e., $|J\cap T|=1$). A lottery then randomly draws a set $L$ of 4 distinct numbers from $S$. A ’J-only Prize’ is awarded if the lottery set $L$ intersects Jen’s set $J$ but is completely disjoint from Tom’s set $T$. The probability of this prize being awarded can be expressed as a fraction $\frac{m}{n}$, where $m$ and $n$ are relatively prime positive integers. Find $m+n$. Analysis: The adapted problem is significantly more difficult and conceptually deeper than the original. The original problem is a standard exercise in combinations and conditional probability. The solution path is straightforward: one simply needs to calculate the number of outcomes for the winning event ("at least two matches") and for the grand prize event ("four matches"), then find their ratio. It primarily tests a student’s fluency with combinatorial formulas. The adapted problem, however, introduces a higher level of complexity by adding a third set, T, with a specific relationship to J ($|J\cap T|=1$). This transforms the problem from a simple counting exercise into one requiring sophisticated logical and set-theoretical reasoning. The key to solving it is not direct calculation, but the conceptual leap of partitioning the universal set S into four distinct subsets based on J and T. Once this structure is understood, the calculation becomes surprisingly simple. In essence, the difficulty shifts from mechanical computation to conceptual abstraction. The adapted problem is harder because it demands a deeper insight into the underlying structure of the sets involved, making it a more elegant and challenging mathematical puzzle.
Figure 16: Task evolution case 9.
Case 10 Original Task: Alice and Bob play the following game. A stack of $n$ tokens lies before them. The players take turns with Alice going first. On each turn, the player removes either 11 token or 44 tokens from the stack. Whoever removes the last token wins. Find the number of positive integers $n$ less than or equal to $2024$ for which there exists a strategy for Bob that guarantees that Bob will win the game regardless of Alice’s play. Evolved Task: An impartial game is played with two stacks of tokens, with sizes $n$ and $m$. Two players take turns making moves. A move consists of removing $k$ tokens from *each* stack. The set of allowed values for $k$ depends on the current minimum stack size, $p=\min(n,m)$: - If $p<20$, the allowed moves are $k\in\{1,2,4\}$. - If $p\geq 20$, the allowed moves are $k\in\{1,4\}$. A move is only possible if both stacks have at least $k$ tokens. The player who makes the last possible move wins. A starting position $(n,m)$ is a ’losing position’ if the second player has a guaranteed winning strategy. Find the total number of losing positions $(n,m)$ such that $1\leq n\leq 100$ and $1\leq m\leq 100$. Analysis: The leap in difficulty from the original to the adapted problem is enormous. The original is a standard one-dimensional game theory exercise where the solution hinges on discovering a simple, repeating pattern in the losing positions (P-positions). It primarily tests pattern recognition. The adapted problem, while appearing to be a more complex two-dimensional game, is a sophisticated test of abstract reasoning. The critical insight, which is far from obvious, is that the game’s two-dimensional state (n, m) collapses. The winning or losing status of any position is determined solely by the minimum of the two stacks, p = min(n, m), and which set of rules applies at that value of p. This reduces the problem to a one-dimensional analysis again, but one that is masked by a misleading setup and complicated by conditional rules. The difficulty is therefore elevated from simple pattern-finding to a much higher level of model simplification and abstraction, followed by a significantly more complex counting phase. It’s a shift from solving a problem to figuring out what the problem really is.
Figure 17: Task evolution case 10.